Part 1- Inventory: You own a toy company and you are producing wooden rocking horses. Assume you have 18,000 units of demand per year. Your ordering cost would be $10 per order. Your unit cost would be $3 and your dollar carrying cost would be 8%
Answer #1


Annual Demand = D = 18000/year

Ordering Cost = Co = $10

Holding Cost = Cc = 8% of $3 = $0.24

Economic Order Quantity Q* = √(2DCo/Cc) = √(2*18000*10*/0.24) = 1224.74

Average Inventory = Q*/2 =  = 1224.74/2 = 612.37

Annual Inventory Holding cost = average inventory * unit carrying cost = (Q*/2)*Cc = (1224.74/2)*0.24 = $146.97

Number of Orders = D/Q* = 18000/1224.74 = 14.7

Annual Order cost = Number of orders * cost/order = (D/Q*)*Co = (18000/1224.74)*10 = $146.97

Total Annual Cost = Annual Inventory Holding Cost + Annual Order Cost = 146.97 + 146.97 = $293.94